# MS-0001-1922.65

### v1

@@ 2.1 Data Collection
We collected 510(k) applicant device submission data and FDA recall data from public FDA datasets related to years 2008 to 2020. The 510(k) applicant device submission data includes the unique 510(k) number of applicant devices, submission and clearance dates as well as device characteristics, such as medical specialty, product type, and device class. Figure 1 demonstrates the recall percentage from 2011 to 2021. The FDA recall data contains the unique 510(k) number of recalled devices, events dates, and severity of recalls. The FDA categorizes recall events into three major classes based on the relative degree of health hazard that can be posed to patients. A Class 1 (severe) recall is a situation with a reasonable probability of serious adverse health consequences or death. A Class 2 (moderate) recall is a situation where temporary or medically reversible adverse health consequences are probable. A Class 3 (mild) recall is a situation that is not likely to cause adverse health consequences.
In the publicly available datasets, there is no direct link between an applicant device and its recall information. For each applicant device, we therefore directly investigated whether it had any prior recalls and identified the recall class when a recall event was reported. Using the public evidence summary documents, we developed a text-scraping algorithm to identify the predicate devices for each cleared 510(k) applicant device. Our resulted data included the applicant devices for which the algorithm identified at least one predicate device. We also extracted device characteristics and recall information for all the predicate devices and added them to our dataset.
@@ 2.2 Variables and Summary
Our primary outcome is a binary recall event indicating whether the applicant device had at least one recall between its FDA clearance date and the end of our study period (2008-2020). A summary of our study sample is provided in Tables 1 and 2.
@@
    Radiology (RA) 3,534 (12.3%) 670 (20.2%)
    Orthopedic (OR) 5,551 (19.4%) 632 (19.0%)
    Cardiovascular (CV) 3,966 (13.9%) 485 (14.6%)
@@
    United States (US) 19,979 (69.8%) 2,699 (81.3%)
    Other 1,887 (6.6%) 179 (5.4%)
@@
    Other_OR 3,180 (11.1%) 368 (11.1%)
    Other_CV 3,016 (10.5%) 346 (10.4%)
    Other_RA 1,495 (5.2%) 238 (7.2%)
    IYN 500 (1.7%) 87 (2.6%)
    JAK 206 (0.7%) 83 (2.5%)
    Other_AN (Anesthesiology) 817 (2.9%) 81 (2.4%)
@@
    0 27914 (97.5%) 3,099 (93.3%)
    1 704 (2.5%) 222 (6.7%)
Note: Data are presented as number (%)
Other = Countries with a frequency of less than 1%
Other_Y = Product codes within medical specialty Y having a frequency of less than 5%
IYN = Ultrasonic pulsed doppler devices; JAK = Radiology Medical diagnostic X-Ray devices
@@
column2 = c, column3 = c, cell221 = c=3, hline1-2,22 = -, & No Recall
(N=28,618)
@@
Num. of Predicate 2.36 (2.29) 2.68 (3.17)
Num. of Unmatched Medical Specialties 0.155 (0.412) 0.153 (0.402)
Num. of Unmatched Product Code 0.507 (0.841) 0.606 (0.987)
Predicate Average Age 4.25 (3.63) 4.04 (3.36)
@@
Predicate Newest Age 3.85 (4.72) 3.34 (4.10)
Predicate Oldest Age 7.44 (6.61) 7.39 (6.59)
@@
Predicate Recalled
    0 23,744 (83.0%) 2,117 (63.7%)
    1 4,874 (17.0%) 1,204 (36.3%)
Number of Predicates Recalled 0.228 (0.615) 0.559 (1.11)
Number of Recalls 0.409 (1.88) 1.28 (4.04)
Variance of Recalls 0.040 (0.895) 0.168 (2.00)
Number of Class 1 0.014 (0.202) 0.035 (0.313)
Number of Class 2 0.387 (1.84) 1.23 (3.97)
Number of Class 3 0.009 (0.103) 0.020 (0.175)
Weighted Number of Recalls 0.247 (1.12) 0.867 (2.57)
Note: Data are presented as number (%), or mean (standard deviation)
Applicant Devices’ Characteristics. The FDA classifies each device into several medical specialties, such as Orthopedic (OR) or Cardiovascular (CV). Medical Specialty refers to the medical specialty of the applicant device identified by the FDA. The FDA classifies medical devices into three classes based on their risks and regulatory controls. Class I devices pose the lowest risk to patients, while Class III devices pose the highest risk. Device Class indicates the FDA device class. Most devices in the 510(k) program fall into Class I or Class II. A limited number of applicant devices (preamendments devices) were initially regulated as Class III with the intent that either the FDA would reclassify the device into a lower class or call for the premarket approval application. Due to a limited number of such applicant devices and the lack of a standard protocol, we only included applicant devices of Class I or II in our analyses. Country Code indicates the country of origin for a device manufacturer. We reduced the number of country code levels by preserving the most common ones and re-coding the others with a frequency of less than 1% as “Other.” The Center for Devices and Radiological Health (CDRH) associates each medical device with a Product Code based upon the medical device function. For example, ultrasonic pulsed doppler devices are assigned to “IYN,” while Ultrasonic Pulsed Echo imaging devices are assigned to “IYO.” We reduced the number of product code levels for each medical specialty by preserving the most common ones and re-coding the others with a frequency of less than 5% as “Other Medical Specialty.” Implantable is a flag indicating whether a device is designed to be placed into a surgically or naturally formed cavity of the human body. Life Sustaining/Supporting is a flag indicating whether a device is essential to restoring or continuing a bodily function.
@@
The first category accounts for the similarity between an applicant device and predicate devices. Number of Predicates in this category is a count of the number of predicate devices identified for an applicant device. Number of Unmatched Medical Specialties measures the unique number of medical specialties for predicate devices, corresponding to an applicant device, that do not match the medical specialty of the applicant device. Similarly, Number of Unmatched Product Codes measures the unique number of product codes for predicate devices, corresponding to an applicant device, that do not match the product code of the applicant device. Table 3 presents the proportion and recall rate of applicant devices where the medical specialty/product code of predicate devices does not match their medical specialty/product code.
Table 3: Proportion and recall rate of applicant devices with predicate devices having mismatched medical specialty/product code
Num. of Unmatched Proportion Recall %
Medical Specialty
0 63.45 60.43
1 26.53 10.54
2 6.90 12.07
3 2.02 13.95
>4 1.10 18.47
Product Code
0 86.05 10.42
1 12.69 10.14
>2 1.26 11.72
@@
The second category accounts for the age of predicate devices. Predicate Average Age is the difference between the average year of approval of predicate devices and the application submission date of the applicant device. Predicate Median Age is the difference between the median year of approval of predicate devices and the application submission date of the applicant device. Predicate Newest Age indicates the difference between the year of approval of the newest predicate device and the application submission date of the applicant device. Similarly. Predicate Oldest Age indicates the difference between the year of approval of the oldest predicate device and the application submission date of the applicant device.
@@
The third category accounts for the recall status of predicates devices. Predicate Recalled is a flag indicating whether at least one of the predicate devices identified for an applicant device is recalled. Number of Predicates Recalled is a count of the number of predicate devices with at least one recall. Number of Recalls is a count of the total number of recalls for the predicate devices. Variance of Recalls is the sample variance of the number of recalls of predicate devices. For example, consider an applicant device with three predicate devices. In the first scenario, the first predicate has six recalls, while the second and third predicates have zero recalls. In the second scenario, each predicate device has two recalls. The sample variance of recalls for the first scenario is 12, while it is zero for the second scenario. We also consider the severity of each recall. Number of Class 1 Recalls is a count of the total number of Class 1 recalls among the predicate devices identified for an applicant device. Number of Class 2 Recalls and Number of Class 3 Recalls can be defined similarly. Another interesting piece of information to consider is the count of recalls with respect to their timing. For example, the importance of a recall event of a predicate device that has occurred many years prior to an applicant device’s submission date may differ from a recent recall event. Weighted Number of Recalls is the weighted number of recalls for the predicate devices identified for an applicant device. We define a time window of ten years such that a recall event is negligible if it has occurred at or more than ten years prior to the applicant’s device submission date. For the other recall events, we assign weights based on the time difference. Thus, a recall event that has occurred at the submission date receives a weight of one, and a recall event that has occurred ten years before the submission date is assigned a weight of zero. For recall events that fall within this ten-year window, weights are calculated proportionally based on their distance from the submission date (e.g., a recall event that happened 4 years ago would be assigned a weight of 0.6).
@@ 3 Machine Learning Models for Predicting Recall Events
We now describe the development and evaluation of ML models for predicting recall events. Our ML models are trained on our data described in the previous section which contains information on a variety of medical devices with different characteristics in order to predict recall events of unseen future applicant devices.
3.1 Machine Learning Models
We consider various ML models and compare their performance to pick the best one. The models we consider are regularized logistic regression using both Lasso and Ridge type penalties, decision tree, random forest, and gradient boosting. The logistic regression model is of the form $\log\left(\frac{\mathbb{P}(Y_{i}=1|X_{i})}{1-\mathbb{P}(Y_{i}=1|X_{i})}\right)=\beta_{0}+\beta_{1}^{\top}X_{i}$, where $X_{i}$ is a vector of information (e.g., characteristics of the applicant device and corresponding predicate devices), $\beta_{1}$ is the unknown coefficient vector, and $\beta_{0}$ is the unknown intercept. Overfitting is one of the biggest causes of the poor performance in working with high-dimensional data such as ours. Regularization is a well-established technique to prevent overfitting by limiting the complexity of the model. The regression coefficients $\beta_{0}$ and $\beta_{1}$ can be estimated by solving the following optimization problem:
$\max_{\beta_{0},\beta_{1}}\left\{\sum_{i=1}^{N}\Big{(}Y_{i}\left(\beta_{0}+\beta_{1}^{\top}X_{i}\right)-\log\left(1+\exp\left(\beta_{0}+\beta_{1}^{\top}X_{i}\right)\right)\Big{)}-\eta\left\lVert\beta_{1}\right\rVert^{p}_{p}\right\},$
where $N$ is the number of observations in the training set, $\left\lVert\cdot\right\rVert_{p}$ denotes $p$-norm, and $\eta$ is the regularization tuning parameter. We implement two types of penalties similar to Lasso and Ridge regression. We do so by setting $p=1$ and $p=2$, respectively. In both models, the value of $\eta$ is tuned through 10-fold cross validation.
We also implement and train a decision tree model, which is better suited to capture nonlinear effects and various interactions in our data. We prune and optimize the depth and the minimum number of samples required to split a node using 10-fold cross validation. In addition, we consider two ensemble learning models built upon Bagging and Boosting techniques. Specifically, we develop a random forest model, which is a popular bagging method obtained by constructing a multitude of de-correlated decision trees. Boosting involves successively training models such that the next model combined with previous ones minimizes the overall prediction error. We also develop a gradient boosting model, which is a popular boosting method obtained by successively constructing different decision trees. Two parameters for random forest (i.e., the number of trees and the maximum depth of each tree) and gradient boosting (i.e., the number of trees and the learning rate) are optimized through 10-fold cross validation.
@@ 3.2 Performance Evaluation of the Machine Learning Models
To compare our ML models, we evaluate their predictive power on unseen 510(k) applicant devices. A 10-fold cross-validation approach is employed to calculate the area under the receiver operating characteristic curve (AUC) for all models. A bootstrapping approach is used to capture statistical fluctuations of AUC scores with respect to random splitting of the data in cross-validation. Table 4 provides a summary of the results, where we report the average AUC for all models, along with the values of standard deviation and quantiles. We find that all models except the decision tree model attain relatively similar performance in terms of the AUC metric, ranging from 0.76 to 0.77. The statistical fluctuations of these four all models are also relatively the same. The lowest standard deviation is 0.027, and the highest 5% and 95% quantiles are 0.72 and 0.81, respectively. The decision tree model is not competitive, which maybe due to overfitting to training sets.
Table 4: Area under the curve for statistical models
@@
Log Reg with Ridge penalty 0.76 0.030 (0.71, 0.80)
Log Reg with Lasso penalty 0.77 0.028 (0.72, 0.81)
Decision tree 0.73 0.036 (0.66, 0.78)
Random forest 0.76 0.031 (0.71, 0.81)
Gradient boosting 0.77 0.027 (0.72, 0.81)
Since gradient boosting has the lowest standard deviation and achieved the highest AUC, we identify it as the best candidate. The receiver operating characteristic (ROC) curve along with the distribution of AUC scores for gradient boosting are depicted in Figure 2. The shaded error bars in Figure 2(a) correspond to ±$\pm 1$ standard deviation, and the vertical lines in Figure 2(b) correspond to different quantiles.
(a) Cross-validation AUC
(b) Distribution of AUC scores
Figure 2: Out-of-sample performance of the selected model (gradient boosting)
@@
The AUC results reported in Table 4 are calculated based on the overall performance of models across all medical specialties. In practice, each medical specialty has a separate committee for evaluating the applicant devices assigned to that medical specialty. The 510(k) process can be viewed as a general rule, but each committee’s evaluation criteria may vary. We evaluate the predictive power of gradient boosting separately for each medical specialty. Figure 3 shows the cross-validation AUC per specialty. As can be seen, the model has the best performance in terms of the AUC metric for applicant devices of Clinical Chemistry (CH), Anesthesiology (AN), and Radiology (RA). On the other hand, Orthopedic (OR), Immunology (IM), and Physical Medicine (PM) are three medical specialties for which the model has the worst performance.
Figure 3: Area under the curve across all specialties
@@
Note: CH = Clinical Chemistry; AN = Anesthesiology; RA = Radiology; HO = General Hospital; NE = Neurology; HE = Hematology; OP = Ophthalmic; GU = Gastroenterology & Urology;
CV = Cardiovascular; TX = Clinical Toxicology; DE = Dental; OB = Obstetrics/Gynecology;
SU = General & Plastic Surgery; MI = Microbiology; EN = Ear, Nose, & Throat; OR = Orthopedic; IM = Immunology; PM = Physical Medicine (PM)
@@
When it comes to a recall event, the severity of the recall can be relativity identified by the FDA’s classification of recalls. Although there is value in detecting each recall class, detecting a high-severity recall class has the highest priority due to its life-threatening nature. We evaluate our gradient boosting model on its ability to detect different classes of recalls. Figure 4 plots the proportion of recalls correctly identified per each recall class versus the overall false recall rate. Overall, the model has better predicting power for a recall Class 1 followed by Class 2 and Class 3. In particular, it achieves the AUC of 0.81, 0.77, and 0.64 in detecting recall Classes 1, 2, and 3, respectively. This is a desired performance, since Class 1 has the highest severity, followed by Class 2 and then Class 3.
@@
Figure 4: Proportion of recall classes correctly identified by our gradient boosting model
@@
Lastly, we investigate the most important variables and their impact when predicating recall risk using our gradient boosting model. We use the shapely additive explanations (SHAP) method (Lundberg and Lee 2017, Lundberg et al. 2020), which leverages a game theory approach to compute the contribution of variables to a predicted value in an additive form. We compute the contribution of each variable to the predicted recall risk in the form of a normalized score ranging between -1 to 1. Figure 5 highlights the 15 most important variables and their impact on the predicted recall risk. They are ordered by decreasing significance. In Figure 5, each point represents a variable’s contribution to the prediction. The value of each contribution is depicted by a color gradient from grey to red, where grey indicates low values and red indicates high values. Our results indicate that Weighted Num. of Recalls and the variables relevant to the age of the predicates are among the top five most significant predictors of the recall risk. Weighted Num. of Recalls accounts for the number of recall events for predicates where the most recent recalls are prioritized. We observe that a higher value of this variable is associated with a higher recall risk. This observation indicates the importance of the timing of predicate recall events, an aspect that has been overlooked in the literature. It also validates our hypothesis that paying attention to predicate devices with recent recalls can go a long way in raising red flags for an applicant device. In particular, a recall event that has happened long ago may have been fully addressed, but some uncertainties might remain unresolved for more recent recalls. Our results also highlight that Num. of Recalls and Variance of Recalls are highly predictive of the recall risk.
@@
With regard to the age of predicates, we observe that a higher value of Predicate Newest Age is associated with a lower risk of recall. Investigating the distribution of the SHAP values corresponding to this variable, we find that devices for which the Predicate Newest Age is less than five years are associated with a significant higher recall risk. This may be because newer predicates might not have undergone extensive usage, which can reveal potential issues over time. We observe that while the Predicate Oldest Age is highly predictive of the recall risk, both high and low values of this variable can lead to a high recall risk. On one hand, older predicates can indicate safety, as they have a history of safe and effective use over time and may be the gold standard for patient care. On the other hand, older predicates may not reflect the advanced technology embedded in new devices, potentially leading to compatibility or performance issues. With regard to the Predicate Average Age and Predicate Median Age, we observe that, overall, lower values of these predictors are associated with a lower risk of recall.

### v2

@@ 2.1 Data Collection
We collected 510(k) applicant device submission data for the years 2008 to 2020 and FDA recall data for the years 2008 to 2021 from public FDA datasets. The 510(k) applicant device submission data includes the unique 510(k) number of applicant devices, submission and clearance dates as well as device characteristics, such as medical specialty, product type, and device class. Figure 1 demonstrates the recall percentage from 2011 to 2021. The FDA recall data contains the unique 510(k) number of recalled devices, event dates, and severity of recalls. The FDA categorizes recall events into three major classes based on the relative degree of health hazard that can be posed to patients. A Class 1 (severe) recall is a situation with a reasonable probability of serious adverse health consequences or death. A Class 2 (moderate) recall is a situation where temporary or medically reversible adverse health consequences are probable. A Class 3 (mild) recall is a situation that is not likely to cause adverse health consequences.
In the publicly available datasets, there is no direct link between an applicant device and its recall information. For each applicant device, we therefore directly investigated whether it had any prior recalls and identified the recall class when a recall event was reported. We developed a text mining algorithm to identify the predicate devices listed by the manufacturers for each cleared 510(k) applicant device using the summary documents associated with each submitted application. The text mining algorithm is described in detail in Appendix 7. Our resulting data included the applicant devices for which the algorithm identified at least one predicate device. Approximately 3% of devices, mainly early scanned summaries, were excluded. The excluded devices are (i) concentrated in 2008–2011, a period that predates FDA’s mandatory eCopy rule (FDA 2020), (ii) spread proportionally across different medical specialties with exclusion rates ranging from 0.9% to 8.2%, and (iii) similarly spread across manufacturing countries, ranging from 0.6% to 5.5%. We also extracted device characteristics and recall information for all the predicate devices and added them to our dataset.
Our dataset, assembled based on the publicly available FDA 510(k) records, contains only those devices that received market authorization. However, FDA does not reveal information on devices that were deemed not substantially equivalent (NSE) and thus rejected, primarily due to confidentiality restrictions (21 CFR 807.95). According to internal analyses from the Center for Devices and Radiological Health (CDRH), approximately 5% of all 510(k) submissions result in NSE decisions (CDRH 2011). While this relatively small fraction of rejections may mitigate concerns about selection bias, it does not eliminate them. In Section 5, we discuss a robustness check conducted to address this limitation in our dataset.
@@ 2.2 Variables and Summary
Our primary outcome is a binary recall event indicating whether the applicant device had at least one recall between its FDA clearance date and the end of our study period. A summary of our study sample is provided in Tables 1 and 2.
@@
    Radiology (RA) 3,532 (12.3%) 664 (20.2%)
    Orthopedic (OR) 5,551 (19.4%) 630 (19.1%)
    Cardiovascular (CV) 3,966 (13.9%) 483 (14.7%)
@@
    United States (US) 19,978 (69.8%) 2,679 (81.4%)
    Other 1,886 (6.6%) 176 (5.3%)
@@
    Other_OR 3,180 (11.1%) 366 (11.1%)
    Other_CV 3,016 (10.5%) 344 (10.4%)
    Other_RA 1,493 (5.2%) 238 (7.2%)
    IYN 500 (1.7%) 86 (2.6%)
    JAK 206 (0.7%) 82 (2.5%)
    Other_AN (Anesthesiology) 817 (2.9%) 79 (2.4%)
@@
    0 27,908 (97.5%) 3,071 (93.3%)
    1 703 (2.5%) 222 (6.7%)
@@
column2 = c, column3 = c, cell221 = c=3, hline1-2,16 = -, & No Recall
(N=28,611)
@@
Num. of Predicate 2.36 (2.29) 2.69 (3.18)
Prop. of Unmatched Medical Specialties 0.077 (0.225) 0.070 (0.212)
Prop. of Unmatched Product Code 0.194 (0.308) 0.210 (0.313)
@@
Predicate Newest Age 3.85 (4.72) 3.32 (4.06)
Predicate Oldest Age 7.44 (6.61) 7.37 (6.55)
@@
Number of Class 1 0.014 (0.202) 0.036 (0.315)
Number of Class 2 0.387 (1.84) 1.22 (3.95)
Number of Class 3 0.009 (0.103) 0.019 (0.175)
Variance of Recalls 0.040 (0.896) 0.163 (1.98)
Weighted Recall Score 0.109 (0.264) 0.262 (0.373)
Note: Data are presented as mean (standard deviation)
The FDA classifies each device into several medical specialties, such as Orthopedic (OR) or Cardiovascular (CV). Medical Specialty refers to the medical specialty of the applicant device identified by the FDA. The FDA classifies medical devices into three classes based on their risks and regulatory controls. Class I devices pose the lowest risk to patients, while Class III devices pose the highest risk. Device Class indicates the FDA device class. Most devices in the 510(k) program fall into Class I or Class II. A limited number of applicant devices (preamendments devices) were initially regulated as Class III with the intent that either the FDA would reclassify the device into a lower class or call for the premarket approval application. Due to a limited number of such applicant devices and the lack of a standard protocol, we only included applicant devices of Class I or II in our analyses. Country Code indicates the country of origin for a device manufacturer. We reduced the number of country code levels by preserving the most common ones and re-coding the others with a frequency of less than 1% as “Other.” The Center for Devices and Radiological Health (CDRH) associates each medical device with a Product Code based upon the medical device function. For example, ultrasonic pulsed doppler devices are assigned to “IYN,” while Ultrasonic Pulsed Echo imaging devices are assigned to “IYO.” We reduced the number of product code levels for each medical specialty by preserving the most common ones and re-coding the others with a frequency of less than 5% as “Other Medical Specialty.” Implantable is a flag indicating whether a device is designed to be placed into a surgically or naturally formed cavity of the human body. Life Sustaining/Supporting is a flag indicating whether a device is essential to restoring or continuing a bodily function.
@@
The first category accounts for the similarity between an applicant device and predicate devices. Number of Predicates in this category is a count of the number of predicate devices identified for an applicant device. Prop. of Unmatched Medical Specialties measures the ratio of the number of distinct medical specialties among an applicant device’s predicates that differ from the applicant device’s own specialty, divided by the total number of predicate devices. Similarly, Prop. of Unmatched Product Codes measures the ratio of the number of distinct product codes among an applicant device’s predicates that differ from the applicant device’s own product code, divided by the total number of predicate devices.
@@
The second category accounts for the age of predicate devices. Predicate Average Age is the difference between the average year of approval of predicate devices and the application submission date of the applicant device. Predicate Median Age is the difference between the median year of approval of predicate devices and the application submission date of the applicant device. Predicate Newest Age indicates the difference between the year of approval of the newest predicate device and the application submission date of the applicant device. Similarly, Predicate Oldest Age indicates the difference between the year of approval of the oldest predicate device and the application submission date of the applicant device.
@@
The third category accounts for the recall status of predicates. Number of Class 1 Recalls is a count of the total number of Class 1 recalls among the predicate devices identified for an applicant device. Number of Class 2 Recalls and Number of Class 3 Recalls can be defined similarly. Variance of Recalls is the sample variance of the number of recalls of predicate devices. For example, consider an applicant device with three predicate devices. In the first scenario, the first predicate has six recalls, while the second and third predicates have zero recalls. In the second scenario, each predicate device has two recalls. The sample variance of recalls for the first scenario is 12, while it is zero for the second scenario. Another interesting piece of information to consider is the timing of the recall events. For example, the importance of a recall event of a predicate device that has occurred many years prior to an applicant device’s submission date may differ from a recent recall event. Weighted Recall Score is calculated as the weighted number of recalls across all predicate devices associated with a given submission, divided by the total number of recalls for the corresponding predicate devices. We define a time window of ten years such that a recall event is negligible if it has occurred at or more than ten years prior to the applicant’s device submission date. For the other recall events, we assign weights based on the time difference. Thus, a recall event that has occurred at the submission date receives a weight of one, and a recall event that has occurred ten years before the submission date is assigned a weight of zero. For recall events that fall within this ten-year window, weights are calculated proportionally based on their distance from the submission date (e.g., a recall event that happened 4 years ago would be assigned a weight of 0.6).
Overall, we constructed 17 predictors for each applicant device, including 11 continuous measures, 3 binary indicator flags (Device Class, Implantable, and Life‑Sustaining/Supporting), and 3 multi‑level categorical variables (Medical Specialty, Product Code, and Country Code). During data preprocessing, we applied one-hot encoding to the multi-level categorical variables, resulting in 20 unique categories for Medical Specialty, 111 for Product Code, and 12 for Country Code.
@@ 3 Machine Learning Models for Predicting Recall Events
We now describe the training and evaluation of ML models for predicting recall events. Our ML models are trained on our data described in the previous section which contains information on a variety of medical devices with different characteristics in order to predict recall events of unseen future applicant devices. We consider various ML models and compare their performance to pick the best one. The models we consider are regularized logistic regression (Log Reg) using both Lasso and Ridge type penalties, decision tree, random forest, and gradient boosting. All models were trained in Python using the scikit-learn (sklearn) library.
To train our ML models and evaluate their performance, we split the data into training (70%) and testing (30%) sets using stratified sampling to preserve class distribution. We then employ a stratified cross-validation approach on the training set to fine-tune hyper-parameters of our ML models and select the best model. Specifically, we perform 10-fold cross-validation on the training set to evaluate the area under the receiver operating characteristic curve (AUC) for various parameter values across all models. This approach helps reduce overfitting and selects a model that generalizes well to unseen data. Using our cross-validation approach, we optimize the regularization parameter for logistic regression models. Similarly, for decision trees, we fine-tune the depth and the minimum number of samples required to split a node. For ensemble methods, we optimize two key parameters for each model: the number of trees and the maximum depth for random forests, and the number of trees and the learning rate for gradient boosting. The implementation details of ML models and cross-validation procedures are outlined in Appendix 8.
@@ 3.1 Performance Evaluation of the Machine Learning Models
To compare our ML models, we evaluate their predictive power using the AUC metric. A bootstrapping approach is used to capture statistical fluctuations of AUC scores with respect to random splitting of the data in cross-validation. Table 3 provides a summary of the results, where we report the average AUC for all models, along with the values of standard deviation and quantiles. We find that all models except the decision tree model attain relatively similar performance in terms of the AUC metric, ranging from 0.77 to 0.78. The statistical fluctuations of these four models are also relatively the same. The lowest standard deviation is 0.022, and the highest 5% and 95% quantiles are 0.74 and 0.81, respectively. The decision tree model underperforms compared to the other models by a slight margin. Based on our further investigation, we find that its relatively weaker performance is not primarily due to overfitting, but is more likely explained by the high predictive power of continuous variables, which is a setting in which single decision trees typically struggle to capture complex relationships.
To further refine our ML models, we explored several common strategies. First, to mitigate class imbalance, we applied the synthetic minority over-sampling technique (SMOTE) to balance the training data. We also used class weighting technique, which adjusts the loss function to penalize misclassification of minority class instances more heavily. However, neither of these strategies led to meaningful improvements in model performance in terms of AUC. Second, since three of our predictors (Medical Specialty, Product Code, and Country Code) contain a large number of unique values, we also tested machine learning models better suited for handling high-cardinality categorical variables in tree-based ML models. Specifically, we used the CatBoost and LightGBM packages, but the trained models did not outperform our models listed in Table 3. Finally, devices cleared by the FDA close to the end of our study period may have had less time to experience a recall than those cleared earlier. This could create a censoring effect for the most recent applicant devices. We employed two robustness checks to assess the potential impact of this censoring. The results suggest that our selected model is not impacted and provides a reliable method of recall risk prediction (see Appendix 9).
Table 3: Area under the curve for ML models
@@
Log Reg with Ridge penalty 0.77 0.022 (0.73, 0.80)
Log Reg with Lasso penalty 0.77 0.022 (0.73, 0.80)
Decision tree 0.75 0.024 (0.71, 0.78)
Random forest 0.77 0.022 (0.73, 0.80)
Gradient boosting 0.78 0.023 (0.74, 0.81)
Among the different models, we ultimately chose gradient boosting for two reasons. First, it consistently showed better performance. Second, its capacity to capture more complex relationships (e.g., potential interactions) makes it more adaptable if the FDA incorporate additional features in the future. The distribution of cross-validation AUC scores along with the out-of-sample ROC curve for gradient boosting are depicted in Figure 4, where the vertical lines in Figure 4 correspond to different quantiles.
Figure 2: Out-of-sample ROC curve
Figure 3: Distribution of cross-validation AUC scores
Figure 4: Performance of the selected model (gradient boosting)
@@
The AUC results reported in Table 3 are calculated based on the overall performance of models across all medical specialties. In practice, each medical specialty has a separate committee for evaluating the applicant devices assigned to that medical specialty. The 510(k) process can be viewed as a general rule, but each committee’s evaluation criteria may vary. We evaluate the predictive power of gradient boosting separately for each medical specialty. Figure 5 shows the out-of-sample AUC per specialty. As can be seen, the model has the best performance in terms of the AUC metric for applicant devices of Clinical Toxicology (TX), Immunology (IM), and Anesthesiology (AN). On the other hand, Hematology (HE), Physical Medicine (PM), and Microbiology (MI) are three medical specialties for which the model has the worst performance.
Figure 5: Out-of-sample AUC across all specialties
Note: TX = Clinical Toxicology; AN = Anesthesiology; IM = Immunology; RA = Radiology;
@@
HO = General Hospital; NE = Neurology; CV = Cardiovascular; GU = Gastroenterology & Urology; CH = Clinical Chemistry; EN = Ear, Nose, & Throat; OB = Obstetrics/Gynecology;
OP = Ophthalmic; DE = Dental; SU = General & Plastic Surgery; OR = Orthopedic;
HE = Hematology; MI = Microbiology; PM = Physical Medicine (PM)
@@
When it comes to a recall event, the severity of the recall can be relatively identified by the FDA’s classification of recalls. Although there is value in detecting each recall class, detecting a high-severity recall class has the highest priority due to its life-threatening nature. We evaluate our gradient boosting model on its ability to detect different classes of recalls. Figure 6 plots the proportion of recalls correctly identified per each recall class versus the overall false recall rate. Overall, the model has better predictive power for Class 1 recalls, followed by Class 2 and Class 3. In particular, it achieves an AUC of 0.78, 0.76, and 0.75 in detecting recall Classes 1, 2, and 3, respectively. This is a desired performance, since Class 1 has the highest severity, followed by Class 2 and then Class 3.
@@
Figure 6: Proportion of recall classes correctly identified versus the overall false recall rate
@@
Lastly, we investigate the most important variables and their impact when predicting recall risk using our gradient boosting model. We use the shapley additive explanations (SHAP) method (Lundberg and Lee 2017, Lundberg et al. 2020), which leverages a game theory approach to compute the contribution of variables to a predicted value in an additive form. We compute the contribution of each variable to the predicted recall risk. Figure 7 highlights the 15 most important variables and their impact on the predicted recall risk. They are ordered by decreasing significance. In Figure 7, each point represents a variable’s contribution to the prediction. The value of each contribution is depicted by a color gradient from grey to red, where grey indicates low values and red indicates high values. Our results indicate that variables related to applicant device characteristics, the age of predicate devices, and the recall status of predicate devices are important predictors of recall risk. Weighted Recall Score accounts for the timing of predicate recall events. We observe that a higher value of this variable is associated with a higher recall risk. We also find that Num. of Class 2 Recalls is highly predictive of the recall risk. These observations not only indicate that the number of predicate recall events matters, but also highlight the importance of their timing, an aspect that has been overlooked in the literature. This supports our hypothesis that paying attention to predicate devices with recent recalls can go a long way in raising red flags for an applicant device. In particular, a recall event that has happened long ago may have been fully addressed, but some uncertainties might remain unresolved for more recent recalls. Our results also highlight that the Variance of Recalls is predictive of the recall risk.
@@
With regard to the age of predicates, we observe that a higher value of Predicate Newest Age is associated with a lower recall risk. This may be because newer predicates might not have undergone extensive usage, which can reveal potential issues over time. We observe that high values of Predicate Oldest Age can lead to a low recall risk. This aligns with our earlier observation that older predicates can indicate safety, as they have a history of safe and effective use over time and may be the gold standard for patient care. With regard to Predicate Median Age, we observe that, overall, a higher value of this predictor is associated with a lower recall risk. Altogether, these findings highlight a learning curve effect. That is, if a predicate device has been on the market only briefly, latent problems may remain undiscovered that may lead to a higher risk of future recalls.